OpenStudy (anonymous):

21 feet 10 feet 23 feet 13 feet Olga is trying to figure out the length of a pool at a hotel that is to be built. A one inch door in the model will become a 3 foot door in reality. If the length of the pool is 7 inches on the model, how long will it be when it is built? 21 feet 10 feet 23 feet 13 feet

OpenStudy (anonymous):

21 feet 10 feet 23 feet 13 feet

OpenStudy (anonymous):

@ikram002p

OpenStudy (ikram002p):

A one inch door in the model will become a 3 foot door in reality. 1 inch \(\rightarrow\) 3 feet 6 inch \(\rightarrow\) ? do cross multiplication

OpenStudy (ikram002p):

? inch= 6 inch × 3feet ? is the same to x and its what u want :)

OpenStudy (anonymous):

18

OpenStudy (anonymous):

@ikram002p

OpenStudy (anonymous):

@iGreen i need u

OpenStudy (igreen):

1 inch → 3 feet 6 inch → 18 feet 18 + 3 = 21. So: 7 inch → 21 feet

OpenStudy (igreen):

@rashional

OpenStudy (ikram002p):

wait sorry i meant 7 *3 :P

OpenStudy (anonymous):

Jaime knows that the area of a triangle can be found by the formula A = 1/2bh. If the area of a triangle is 26 m2 and the height is 13 meters, what is the length of the base? 2 meters 8 meters 4 meters 1 meter

OpenStudy (anonymous):

Solve for q if 8/q = 64/16 4 2 8 16

OpenStudy (anonymous):

these the last two @iGreen

OpenStudy (igreen):

Jaime knows that the area of a triangle can be found by the formula A = 1/2bh. If the area of a triangle is 26 m2 and the height is 13 meters, what is the length of the base? Plug what you know in the formula: A = 1/2bh 26 = 1/2b(13) Simplify: Plug in one of the answers in: 26 = 1/2(2)(13) 26 = (1)(13) 26 = 13 WRONG 26 = 1/2(4)(13) 26 = (2)(13) 26 = 26 CORRECT Answer = 4

OpenStudy (igreen):

Solve for q if 8/q = 64/16 64 / 8 = 8 16 / 8 = 2 2 is your answer.

OpenStudy (igreen):

Good luck!